OpenStudy (anonymous):

find the difference: 11/w minus 5/w

OpenStudy (anonymous):

can you explain how to do it please?

OpenStudy (anonymous):

\[\frac{ 11 }{ w }-\frac{ 5 }{ w }=\frac{ 11-5 }{ w }\]

OpenStudy (paki):

11/w-5/w... take lcm and solve...

OpenStudy (paki):

lcm will be 11-5/w = 6/w...

OpenStudy (anonymous):

thank you that makes sense